CLIMM(1) CLIMM(1) NAME climm−asmall,textmodeICQclient SYNOPSIS climm[-h][-c][-b<basedir>][-i<locale>][-v[level]][-l<logplace>][-u<UIN>][-p<passwd>] [-s<status>][-C<commands>] DESCRIPTION climmisatextbasedICQclientsupportingmanyoftheICQfeaturesprovidedbytheclientby Mirabilis.climmallowstochangetheuserpassword,createnewaccounts,updatinguserinfo, searchotherusers,sendSMS,opendirectconnections,sendacknowledgedmessages,transferfiles, detectotheruser’sclient,sendUTF-8encodedmessages,executearbitrarycommandsuponany event,etc. OPTIONS −h|−−help Displayashorthelptextandexit. −c|−−nocolor Disabletheuseofcolor. −b|−−basedir<basedir> Specifybasedirasthebasedirectorywhereclimmwilllookforitsfiles,inparticularits climmrcfile. −i|−−i18n<locale> Usethelocalelocaleinsteadoftheonederivedfromenvironmentvariables. −v|−−verbose[level] Setverbosemode,orsetverbositytolevel. −l|−−logplace<file>|<dir> Usefileasfiletowriteloginto,orusedirasadirectorytocreateperuserlogfilesin. −u|−−uin<uin> Loginwithuinonly.Mayberepeated. −p|−−passwd<passwd> Loginwithpasswdaspassword.Mayberepeatedforeach-uswitch,orsingle. −s|−−status<status> Loginwithstatusasstatus,whichmaybegivennumericallyorinv,dnd,occ,na,away, ffc,oroff[line]. −C|−−cmd<commands> Executeclimmcommandscmdsafterlogininsteadofthedefaultegcommand. USAGE Uponfirststart(forthegivenbasedirectory),climmwillaskforaUIN,orcreateanewone. Afterthefirstlogin,theserverbasedcontactlistwillbeimportedautomatically. Basiccommandsincludemsgtomessageothercontacts(trypressingthe<tab>keyseveraltimes onanemptyinputlinetobringupmsgcommandstomessagecontactsrecentlytalkedto),addto addnewcontacts,update,other,andabouttochangeyourmetadataontheICQserver,and,of course,helpforonlinehelp.Acompletelistofcommandsisavailableintheclimmcmds(7)man page. FEATURES climmisveryportableandrunsonLinux,DOS,OS/2,Win32,BeOS,AmigaOS,GNUHURD andanyothermodernunixsystem.climmversion0.5isknowntorunonLinux,AIX,HPUX, Tru64,AmigaOS,partiallyBeOSandotherunices.However,theotherportsmayormaynotstill workandanyonewishingtotestthemiswelcometo. climm 1 CLIMM(1) CLIMM(1) HISTORY 16Aug1998 MatthewD.Smithreleasesclimmuntotheworld.Followedby12datedreleasesuntil25 Sept1998. 05Jan2001 MatthewD.Smithdiedinacaraccident.UserspickupwhereMatthewleftoff,formalizeclimmundertheBSDlicenseandrestartthedevelopmentofclimm. 10June2002 Version0.4.9isreleasedundertheGPLlicence,finallyimplementingthenewversion8of theICQprotocol,donebyRu¨ digerKuhlmann. MAILING-LIST Theclimmmailinglistisclimm-list@climm.org.Informationaboutsubscribingandarchivesis availableathttp://www.climm.org/pipermail/climm-list/. DOWNLOAD ViaHTTPfromhttp://www.climm.org/download. TogetthenewestversionfromCVS,enter: cvs−d:pserver:anonymous@climm.org:/var/lib/cvscoclimm SEEALSO climmrc(5),climmcmds(7) FILES ˜ /.climmrc Oldconfigurationfileforclimm−obsolete. ˜ /.climm/ Thedefaultbasedirectorywhereclimmlooksforitsfiles. BASE/climmrc Configurationfileforclimm.Containsglobalconfiguration. BASE/status Configurationfileforclimm.Containscontactslistandcontactoptions. BASE/i18n/*.i18n Localoverridesforclimm’stranslationfiles. ˜ /.climm/i18n/*.i18n Localoverridesforclimm’stranslationfiles. BASE/contacts/ Directorywheremetadataforcontactsisstored. BASE/history/ Defaultdirectoryforlogfiles. BASE/files/ Directorywhereincomingfiletransfersaresaved. BASE/remote-control OlddefaultlocationoftheFIFOusedforscripting.Renamedasitconfusedlusers. BASE/scripting DefaultlocationoftheFIFOusedforscripting. PREFIX/share/climm/*.i18n Internationalizationfilescontainingtranslationsofallstringsclimmoutputs. BUGS SendbugreportsorcommentstoRu¨ digerKuhlmann<climm@ruediger-kuhlmann.de>(inGerman orEnglish). climm 2 CLIMM(1) CLIMM(1) AUTHOR ThismanualpagewasderivedbyJamesMorrison<ja2morrison@student.math.uwaterloo.ca>from theminimalmanpagewrittenbyJordiMallach<jordi@sindominio.net>fortheDebian GNU/Linuxsystemandlaterkeptup-to-dateby<climm@ruediger-kuhlmann.de>. climm 3